# TEG-DB: A Comprehensive Dataset and Benchmark of Textual-Edge Graphs

### v1

@@ 5.2 Effectiveness Analysis for Link Prediction
In this subsection, we analyze the link prediction from the various models applied in the study. Table 3 and 3 represent the effect of link prediction on different datasets from various distinct. The results on other datasets can be found in Appendix B.1. We can further draw several observations from Table 3 and 3. First, For PLM-based and GNN-based methods, the state-of-the-art methods for Goodreads-Children and Goodreads-Crime datasets are GeneralConv and GINE, respectively. Under the condition of using the same embeddings, they outperform the worst method by approximately 5% and 7% in terms of AUC and F1 across these two datasets. For the Amazon-Apps and Amazon-Movie datasets, the state-of-the-art methods are EdgeConv and GeneralConv. They outperform the worst method by approximately 3% and 7% in terms of AUC and F1 for Amazon-Apps, and by 8% and 7% in terms of AUC and F1 for Amazon-Movie, respectively. For the Citation and Twitter datasets, the state-of-the-art method is GraphTransformer. It outperforms the worst method by approximately 20% and 30% in terms of AUC and F1 for Citation, and by 12% and 9% in terms of AUC and F1 for Twitter, respectively. Second, For the LLM as Predictor methods, we find that they do not perform well in predicting links. The best method among them has an AUC and F1 gap of approximately 10% - 30% compared to the best PLM-based and GNN-based methods for all datasets. Third, Using edge text provides at least approximately a 3% improvement in AUC and at least approximately an 8% improvement in F1 compared to not using edge text for all datasets.
@@ 5.3 Effectiveness Analysis for Node Classification
In this subsection, we analyze the node classification results from various models. Table 5 and 5 display the impact on different datasets from various distinct, with additional results in Appendix B.2. We can derive some insights from the data. First, for PLM-based and GNN-based methods, the state-of-the-art models for Goodreads-Children and Goodreads-Crime are GraphSAGE and GeneralConv, respectively, outperforming the worst method by approximately 8% and 20% in AUC-micro and F1-micro for Goodreads-Children, and by 4% and 15% for Goodreads-Crime. In the E-commerce domain, GraphSAGE is the top method for Amazon-Apps and Amazon-Movie, outperforming the worst method by about 10% and 6% in AUC-micro and F1-micro for Amazon-Apps, and by 1% and 10% for Amazon-Movie. GINE and EdgeConv also show superior performance, exceeding the worst method by approximately 35% and 40% in ACC and F1 for Citation, and by 5% and 12% for Twitter. Second, LLM as Predictor methods perform poorly in node classification, with the best method showing an AUC-micro gap of about 30% compared to the best PLM-based and GNN-based methods. Their low F1-micro score could be due to the large number of predicted categories. Third, incorporating edge text results in at least a 3% improvement in AUC-micro and a 6% improvement in F1-micro across all datasets, compared to not using edge text.
Observation. (1) The state-of-the-art model varies across different datasets. Data variability and complexity play significant roles in influencing model performance. (2) Edge text is crucial for TEG tasks. Including edge text enriches relationship information, enabling a more precise depiction of interactions and relationships between nodes, which enhances overall model performance. (3) The scale of PLMs significantly impacts the performance of TEG tasks. Larger model scales result in higher-quality text embeddings and better semantic understanding, leading to improved model performance. (4) When using LLMs as predictors, they struggle to fully comprehend graph topology information. LLMs are designed for linear sequence data and do not inherently capture the complex relationships and structures present in graph data, leading to lower performance on TEGs link prediction and node classification.
@@ 6 Discussion
Textual-Edge graphs have emerged as a prominent graph format, which finds extensive applications in modeling real-world tasks. Our research focuses on comprehensively understanding the textual attributes of nodes and their topological connections. Furthermore, we believe that exploring strategies to enhance the efficiency of LLMs in processing TEGs is deemed meaningful. Despite the proven effectiveness of LLMs, their operational efficiency, especially in managing TEGs, poses a significant challenge. Notably, employing APIs like GPT4 for extensive graph tasks may result in considerable expenses under current billing models. Additionally, deploying open-source large models such as LLaMa for tasks like parameter updates or inference in local environments demands substantial computational resources and storage capacity. Moreover, the constraints imposed by context windows in LLMs also impact their effectiveness in encoding node and edge text within TEGs.
@@ A.1 Dataset format
For each dataset, all unprocessed raw files are represented in .json format. After preprocessing, we store the graph-type data compatible with PyTorch Geometric (PyG) [9] in the .pt format using PyTorch. Specifically, we have retained the raw text on nodes, the labels on nodes, the raw text on edges, and the adjacency matrix. We uniformly store the text embeddings of node and edge text in .npy files and load them during data processing.
@@ Appendix B Experiment
B.1 Effectiveness Analysis for Link Prediction
@@
B.2 Effectiveness Analysis for Node Classification

### v3

@@ 5.2 Effectiveness Analysis for Link Prediction
In this subsection, we analyze the link prediction from the various models applied in the study. Table 2 and 3 represent the effect of link prediction on different datasets from various distinct models. The results on other datasets can be found in Appendix B.2. We can further draw several observations from Table 2 and 3. First, For PLM-based and GNN-based methods, the state-of-the-art methods for Goodreads-Children and Goodreads-Crime datasets are both GeneralConv. Under the condition of using the same embeddings, they outperform the worst method by approximately 5% and 7% in terms of AUC and F1 across these two datasets. For the Amazon-Apps and Amazon-Movie datasets, the state-of-the-art methods are EdgeGNN and GeneralConv. They outperform the worst method by approximately 3% and 7% in terms of AUC and F1 for Amazon-Apps, and by 8% and 7% in terms of AUC and F1 for Amazon-Movie, respectively. For the Citation and Twitter datasets, the state-of-the-art method is GraphTransformer. It outperforms the worst method by approximately 20% and 30% in terms of AUC and F1 for Citation, and by 12% and 9% in terms of AUC and F1 for Twitter, respectively. Second, Entangled-GPT methods, which entangle edge text and node text first before encoding with GPT consistently outperform the approach of directly encoding the text through GPT, yielding about 2% improvement in both AUC and F1 metrics across all datasets on the link prediction tasks. Third, For the LLM as Predictor methods, we find that they do not perform well in predicting links. The best method among them has an AUC and F1 gap of approximately 10% - 30% compared to the best PLM-based and GNN-based methods for all datasets. Fourth, using edge text provides at least approximately a 3% improvement in AUC and at least approximately an 8% improvement in F1 compared to not using edge text for all datasets.
@@ 5.3 Effectiveness Analysis for Node Classification
In this subsection, we analyze the node classification results from various models. Table 4 and 5 display the impact on different datasets from various distinct, with additional results in Appendix B.3. We can derive some insights from the data. First, for PLM-based and GNN-based methods, the state-of-the-art models for Goodreads-Children and Goodreads-Crime are GraphSAGE and GeneralConv, respectively, outperforming the worst method by approximately 8% and 20% in AUC-micro and F1-micro for Goodreads-Children, and by 4% and 15% for Goodreads-Crime. In the E-commerce domain, GraphSAGE is the top method for Amazon-Apps and Amazon-Movie, outperforming the worst method by about 10% and 6% in AUC-micro and F1-micro for Amazon-Apps, and by 1% and 10% for Amazon-Movie. GINE and EdgeConv also show superior performance, exceeding the worst method by approximately 35% and 40% in ACC and F1 for Citation, and by 5% and 12% for Twitter. Second, Entangled-GPT methods outperform the approach of directly encoding the text through GPT, yielding about 2% improvement in both AUC and F1 metrics across all datasets on the node classification tasks. Third, LLM as Predictor methods perform poorly in node classification, with the best method showing an AUC-micro gap of about 30% compared to the best PLM-based and GNN-based methods. Their low F1-micro score could be due to the large number of predicted categories. Third, incorporating edge text results in at least a 3% improvement in AUC-micro and a 6% improvement in F1-micro across all datasets, compared to not using edge text.
Observation. (1) The state-of-the-art model varies across different datasets. Data variability and complexity play significant roles in influencing model performance. (2) Edge text is crucial for TEG tasks. Including edge text enriches relationship information, enabling a more precise depiction of interactions and relationships between nodes, which enhances overall model performance. (3) Encoding text in an entangled manner is more beneficial for avoiding information loss. The advantage of this method over existing approaches is its ability to effectively preserve the semantic relationships between nodes and edges, making it more suitable for capturing complex relationships. (4) The scale of PLMs significantly impacts the performance of TEG tasks, especially on datasets with rich text on nodes and edges. Larger model scales result in higher-quality text embeddings and better semantic understanding, leading to improved model performance. (5) When using LLMs as predictors, they struggle to fully comprehend graph topology information. LLMs are designed for linear sequence data and do not inherently capture the complex relationships and structures present in graph data, leading to lower performance on TEGs link prediction and node classification.
@@ 6 Discussion
Textual-Edge graphs have emerged as a prominent graph format, which finds extensive applications in modeling real-world tasks. Our research focuses on comprehensively understanding the textual attributes of nodes and their topological connections. Furthermore, we believe that exploring strategies to enhance the efficiency of LLMs in processing TEGs is deemed meaningful. Despite the proven effectiveness of LLMs, their operational efficiency, especially in managing TEGs, poses a significant challenge. Notably, employing APIs like GPT4 for extensive graph tasks may result in considerable expenses under current billing models. Additionally, deploying open-source large models such as LLaMa for tasks like parameter updates or inference in local environments demands substantial computational resources and storage capacity. Please refer to the Appendix C for more details.
@@ 7 Conclusion
Checklist
The checklist follows the references. Please read the checklist guidelines carefully for information on how to answer these questions. For each question, change the default [TODO] to [Yes] , [No] , or [N/A] . You are strongly encouraged to include a justification to your answer, either by referencing the appropriate section of your paper or providing a brief inline description. For example:
  • Did you include the license to the code and datasets? [Yes] See Section A.2.
  • Did you include the license to the code and datasets? [No] The code and the data are proprietary.
  • Did you include the license to the code and datasets? [N/A]
Please do not modify the questions and only use the provided macros for your answers. Note that the Checklist section does not count towards the page limit. In your paper, please delete this instructions block and only keep the Checklist section heading above along with the questions/answers below.
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
      Answer:[Yes]
      Justification: See Section abstract and introduction.
    2. (b)
      Did you describe the limitations of your work?
      Answer:[Yes] See Appendix D
    3. (c)
      Did you discuss any potential negative societal impacts of your work?
      Answer:[N/A]
      Justification: Our paper has no potential negative social impacts.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them?
      Answer:[Yes]
      Justification: Our paper strongly conforms to the ethics review guidelines.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results?
      Answer:[N/A]
      Justification: No results requiring assumptions.
    2. (b)
      Did you include complete proofs of all theoretical results?
      Answer:[N/A]
      Justification: No theoretical results requiring proof.
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)?
      Answer:[Yes]
      Justification: We include a URL in the abstract.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)?
      Answer:[Yes]
      Justification: See Section 5.1 .
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)?
      Answer:[Yes]
      Justification: See Appendix 5.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)?
      Answer:[Yes]
      Justification: See Appendix 5.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators?
      Answer:[Yes]
    2. (b)
      Did you mention the license of the assets?
      Answer:[Yes]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL?
      Answer:[Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating?
      Answer:[Yes]
      Justification: See Section 4.2.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content?
      Answer:[N/A]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable?
      Answer:[N/A]
      Justification: This work does not involve crowdsourcing or research with human subjects.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable?
      Answer:[N/A]
      Justification: This work does not involve crowdsourcing or research with human subjects.
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation?
      Answer:[N/A]
      Justification: This work does not involve crowdsourcing or research with human subjects.
@@ A.1 Dataset format
For each dataset, all unprocessed raw files are represented in .json format. After preprocessing, we store the graph-type data compatible with PyTorch Geometric (PyG) Fey and Lenssen [2019] in the .pt format using PyTorch. Specifically, we have retained the raw text on nodes, the labels on nodes, the raw text on edges, and the adjacency matrix. We uniformly store the text embeddings of node and edge text in .npy files and load them during data processing.
@@ Appendix B Experiment
B.1 Implementation Details
GNNs are mainly derived from the implementation in the PyG library Fey and Lenssen [2019]. For the node classification task, numerical node labels corresponding to the nodes within the graph are necessary. This involves converting the categorical node categories found in the original data into numerical node labels within the graph. For the link prediction, we randomly sample node pairs that do not exist in the graph as negative samples, along with some edges present as positive samples. For LLM-based predictor methods, we focus on node classification and link prediction tasks. For node classification, inspired by the recent LLM-based classification algorithm Liu et al. [2024], we use GPT-4 and GPT-3.5-TURBO models to predict the classification of text nodes by providing the probability for each class. We randomly select 1,000 text nodes along with all classification labels for this task. For the link prediction task, we also apply the GPT-4 and GPT-3.5-TURBO models to determine whether two text edges are related, providing an answer with the corresponding probability. For this task, we randomly select 1,000 pairs of positive text edge indices from the graph and an equal number of negative edges.
B.2 Effectiveness Analysis for Link Prediction
@@
B.3 Effectiveness Analysis for Node Classification
@@
Appendix C Discussion
Notably, employing APIs like GPT4 for extensive graph tasks may result in considerable expenses under current billing models. Additionally, deploying open-source large models such as LLaMa for tasks like parameter updates or inference in local environments demands substantial computational resources and storage capacity. Consequently, enhancing the efficiency of LLMs for graph-related tasks remains a critical concern. Moreover, the constraints imposed by context windows in LLMs also impact their effectiveness in encoding node and edge text within TEGs.
Appendix D Limitation
Comprehensive evaluation of tasks often demands significant computational resources, which can be a burden for researchers and smaller organizations.